# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

### v2

@@ VI Conclusion
This study has presented novel controller design strategy to achieve iso-damping robust control. The proposed approach is simpler, more practical, and reliable than conventional iso-damping controller design methodologies. The proposed approach designs the iso-damping controller on the basis of the MR-D3C framework, requiring only one-shot input/output data. The plant model, the computation of the frequency response, or repeated experiments are unnecessary. The present controller tuning scheme explicitly evaluates the BIBO stability of the resultant closed-loop system from a reference input to an output. The numerical examples have demonstrated the effectiveness of the proposed approach. Moreover, this study has verified the advantage of FO control for iso-damping robust control.
In the future, we will analyze the stability and performance using the present approach in a finite sample under measurement noise.

### v3

@@ VI Conclusion
This study has presented novel controller design strategy to achieve iso-damping robust control. The proposed approach is simpler, more practical, and reliable than conventional iso-damping controller design methodologies. The proposed approach designs the iso-damping controller on the basis of the MR-D3C framework, requiring only one-shot input/output data. The plant model, the computation of the frequency response, or repeated experiments are unnecessary. The present controller tuning scheme explicitly evaluates the BIBO stability of the resultant closed-loop system from a reference input to an output. The numerical and experimental studies have demonstrated the effectiveness of the proposed approach. Moreover, this study has verified the advantage of FO control for iso-damping robust control.
In the future, we will analyze the stability and performance of the present approach with a finite data length under measurement noise.